# Enhancing Fault Detection for Large Language Models via Mutation-Based Confidence Smoothing

### v1

@@ Abstract.
Large language models (LLMs) achieved great success in multiple application domains and attracted huge attention from different research communities recently. Unfortunately, even for the best LLM, there still exist many faults that LLM cannot correctly predict. Such faults will harm the usability of LLMs. How to quickly reveal them in LLMs is important, but challenging. The reasons are twofold, 1) the heavy labeling effort for preparing the test data, and 2) accessing closed-source LLMs such as GPT4 is money-required. To handle this problem, in the traditional deep learning testing field, test selection methods have been proposed for efficiently testing deep learning models by prioritizing faults. However, the usefulness of these methods on LLMs is unclear and under exploration. In this paper, we first study the effectiveness of existing fault detection methods for LLMs. Experimental results on four different tasks (including both code tasks and natural language processing tasks) and four LLMs (e.g., LLaMA and GPT4) demonstrated that existing fault detection methods cannot perform well on LLMs (e.g., seven out of eight methods perform worse than random selection on LLaMA). To enhance existing fault detection methods, we propose MuCS, a prompt Mutation-based prediction Confidence Smoothing method for LLMs. Concretely, we mutate the prompts and compute the average prediction confidence of all mutants as the input of fault detection methods. The results show that our proposed solution significantly enhances existing methods with the improvement of test relative coverage by up to 97.64%.
@@ 1. Introduction
Since the last few years, the potential of large language models (LLMs) has brought hope to general artificial intelligence (AGI). LLMs achieved promising and even the best results compared to other methods in many areas, such as question answering (Zhou et al., 2022), sentiment analysis (Zhang et al., 2023a), and code understanding (Ma et al., 2023) that make LLMs become the first choice of deep learning models in such tasks. Besides the most famous application ChatGPT 111OpenAI, ChatGPT, https://chat.openai.com/, multiple LLMs have been proposed, e.g., LLaMA (Touvron et al., 2023), Alpaca (Taori et al., 2023), and Cerebras-GPT (Dey et al., 2023). More interestingly, focusing on software engineering (SE) tasks, a series of code-related LLMs with good programming ability have been released to help the software development such as StarCoder (Li et al., 2023a) and CodeLLaMA (Roziere et al., 2023). Learning to use LLMs is already a basic skill for us in daily life.
@@
Despite the success, the same as other types of deep learning models, LLMs also suffer from different problems that limit their practical usage. For example, LLMs sometimes generate unrelated outputs to the inputs or incorrect results misaligned with established world knowledge, which is known as the hallucination problem (Zhang et al., 2023b). Besides, recent work (Dong et al., 2023a) showed that LLMs can be easily fooled by adversarial attacks and are not robust. Furthermore, considering SE tasks, existing code-related LLMs cannot capture equivalent semantics when the given natural language prompt expresses the same meaning in different languages (Peng et al., 2024). Those limitations remind us it is necessary to comprehensively evaluate and test the capabilities of LLMs before deployment.
@@
The basic way to test LLMs is to prepare as much as possible test data to assess the performance of LLMs accordingly. However, preparing such test data is heavy work due to the complicated process of data collection, data cleaning, and data labeling. Especially, data labeling requires human effort with domain knowledge which is the most challenging part. Moreover, some LLMs are closed-source which requires funding expenses, e.g., it tasks $0.03/1K tokens of inputs for GPT4 using OpenAI API. Therefore, it is almost impossible to test LLMs for people who have limited budgets. How to efficiently test the performance of LLMs with less effort becomes an urgent problem.
@@
To tackle the data labeling issue, in the field of deep learning testing, multiple fault detection methods that quickly identify fault data (data the model has incorrect prediction) in the test set without using the labeling information (Hu et al., 2024) have been proposed. Based on the information required, existing methods can be divided into learning-based methods and output-based methods. Those methods have been widely studied in the classical deep learning models and have proven to be effective in saving the labeling budget during testing. However, there are no works that explore the usefulness of fault detection methods for LLMs. It is unclear whether we can directly employ such methods to test LLMs and save our effort or not.
@@
To bridge this gap, in this paper, we first study the effectiveness of existing fault detection methods on LLMs. The major challenge in conducting such a study is that most existing fault detection methods are proposed for classification tasks and use the output probabilities for the selection. It is necessary to design the prompt to guide LLMs to produce the output with their confidence. To solve this, for datasets that have more than two classes, we add examples in each class in the prompt as guidance to ask LLMs to predict new data. In total, our study covers nine fault detection methods including both learning-based methods, e.g., TestRank (LI et al., 2021), and output-based methods, e.g., ATS (Gao et al., 2022). Based on the experiment results collected from four datasets (e.g., sentiment analysis and code clone detection) and four LLMs (e.g., LLaMA and GPT4), we found that 1) GPT models are more confident in their outputs (e.g., 0.84 vs. 0.76 of GPT3.5 vs. CodeLLaMA), 2) LLMs are not well-calibrated and overconfident in clone detection, problem classification, and news classification tasks, and 3) existing fault detection methods perform pool on LLMs compared their performance on classical deep learning models. For example, seven out of eight methods perform worse than random selection on LLaMA. There is a need to propose methods to further enhance existing fault detection methods.
@@
To do so, inspired by (Li et al., 2023b; Wang et al., 2021) which utilize mutation testing techniques to help test deep learning models, we propose MuCS, a prompt Mutation-based prediction Confidence Smoothing method to enhance existing fault detection methods. Specifically, we mutate the prompt by transformation methods to generate a set of prompt mutants first. Here, both text augmentation methods and code refactoring methods have been considered in MuCS. Then, we collect the output probabilities of all mutants and compute the average prediction confidence. Finally, we use the averaged confidence to perform fault detection using current fault detection methods. We compare the effectiveness of fault detection methods with and without our confidence smoothing method and found that using MuCS significantly enhances the performance of existing methods by up to 97.62%.
@@
This is the first study that explores the effectiveness of fault detection methods on LLMs. We found that 1) existing methods that perform well on traditional deep learning models cannot correctly detect faults in LLMs; 2) LLMs are overconfident in clone detection, problem classification, and news classification tasks; 3) the prediction confidence of LLMs is concentrated in a few intervals.
@@ 2.1. Fault Detection for Deep Neural Networks
The early method DeepGini (Feng et al., 2020) has been proposed by Feng et al. to reveal inputs that are more likely been mispredicted by the model. DeepGini defined a new Gini score to measure the uncertainty of deep learning models on the inputs. Later on, inspired by the mutation testing in the conventional SE field, Chen et al. proposed to mutate both input samples and deep learning models and identify faults by computing the killing score. Their evaluation on more than 20 tasks demonstrated that mutation testing is useful for helping find bugs in deep learning models. Furthermore, Gao et al. proposed an adaptive test selection (ATS) method. ATS not only considers detecting faults but also tries to find diverse faults (diverse here means faults are from different categories) in deep learning models. More recently, Bao et al. first empirically proved that simple methods (i.e., output probability-based methods) perform well on fault detection. Then, they proposed a new method to compute the uncertainty scores by averaging the basic score of the input and scores from the neighbors of this input. Besides, Ma et al. conducted a comprehensive empirical study to explore the potential of test selection methods and active learning methods on fault detection. They found that MaxP is the best in terms of the correlation between the MaxP score and the misclassification. Besides the SE community, researchers from the ML community also contributed to this field. Dan et al. built the first benchmark for detecting misclassified data and found that just using the confidence score is already a strong baseline for this task. Li et al. proposed TestRank which uses graph neural networks to learn the difference between benign inputs and faults for fault detection.
Different from these works, our work is the first to explore the effectiveness of fault detection methods for large language models. We found that existing methods cannot well detect faults in LLMs and proposed the use of prompt mutation to further enhance these methods.
@@ 2.2. Large Language Model Testing
Large language models have been the most used deep learning models for most of the tasks. Comprehensively testing or evaluating the performance of LLMs becomes one of the most important directions in all the communities.
@@
More recently, Yuan et al. (Yuan et al., 2023) compared LLMs with conventional code models that are fine-tuned for specific code tasks. They found that in the zero-shot setting, instruction-tuned LLMs have competitive performance to fine-tuned code models. In the one-shot setting, the guidance by one-shot example sometimes harms the performance of LLMs. Du et al. (Du et al., 2024) argued that existing works that evaluate the effectiveness of LLMs for code mainly focus on simple code tasks (function-level code synthesis) and built the first class-level cod regeneration benchmark for the LLM evaluation. Based on their benchmark, they found that existing LLMs have significantly lower performance on class-level code generation tasks compared to method-level code generation tasks. There is still a big room to improve the ability of LLMs for code generation.
Unlike existing works that focus on evaluating the performance of LLMs from the perspective of preparing test cases, our work measures the capability of LLMs from another view and studies the problem of quickly identifying potential faults in LLMs to save the testing cost.
@@ 3. Empirical Study
In this work, we only focus on the problem of fault detection of LLMs. We consider both learning-based fault detection methods and output-probability-based methods in the literature. Based on the previous definition (Hu et al., 2024), given a LLM , an unlabeled test set , and a labeling budget , fault detection is to select a subset of such that where are the labels corresponding to , and is the performance measurement function. We follow the guidance of previous works (Hu et al., 2024; Ma et al., 2021; Hu et al., 2023) to select nine fault detection methods in our study. We exclude neuron coverage-based methods and surprise adequacy methods in our benchmark since the parameter size of LLMs is too big (>7B) and the coverage and adequacy score are difficult to compute.
@@ 3.1. Fault Detection Methods
Let be a set of test inputs and be the corresponding label set, and denote a given test sample and its true label, respectively. Let be the likelihood of belonging to class produced by the model .
@@
Max Probability (MaxP) prioritizes data based on the maximum output probability . The data sample with a lower MaxP score is more likely to be a fault.
@@ 3.2. Research Question
Our study plans to explore the effectiveness of existing fault detection methods on LLMs. Concretely, we answer the following two research questions:
RQ1: How confident are LLMs in their prediction? Most (seven out of nine) of the fault detection methods are purely output uncertainty-based. Therefore, checking the prediction confidence produced by LLMs before running fault detection methods is necessary.
@@
3.3. Datasets and Models
Task | Class Number | Test Size | Model | Accuracy |
CodeLLaMA | 49.0% | |||
GPT3.5 | 75.5% | |||
Clone Detection (CD) | 2 | 200 | GPT4 | 77.0% |
CodeLLaMA | 17% | |||
GPT3.5 | 36% | |||
Problem Classification (PC) | 5 | 200 | GPT4 | 100.0% |
LLaMA | 89.3% | |||
GPT3.5 | 82.0% | |||
Sentiment Analysis (Sentiment) | 3 | 150 | GPT4 | 90.0% |
LLaMA | 21.5% | |||
GPT3.5 | 60.0% | |||
TagMyNews (TMN) | 7 | 200 | GPT4 | 82.0% |
In total, we consider four datasets including both natural language classification tasks and programming language classification tasks, and four types of LLMs covering both open-sourced models and closed-source models.
Clone Detection (CD): Utilizing the BigCloneBench (Svajlenko et al., 2014) dataset, we randomly selected 200 items outside of 2 classes to investigate code clone detection. Models are evaluated on their ability to predict the presence of code clones (a probability score from 0 to 1, where 0 indicates no clone and 1 indicates a clone) between code snippet pairs, a critical task for enhancing software maintenance and development practices.
Problem Classification (PC): Based on the Java250 (Puri et al., 2021) dataset from the CodeNet Project, problem classification focuses on the classification of programming challenges on LLMs. A set of 200 samples across 5 distinct problem types was chosen in a way of a one-shot learning prompt. This task tests the model’s proficiency in categorizing coding problems using a singular example for each category, requiring output as a probability distribution across classes that sum to 1.
Sentiment Analysis (Sentiment): For this task, 150 samples were extracted from the Sentiment Analysis of IMDB Movie Reviews dataset 222Autolabel, https://github.com/refuel-ai/autolabel. The challenge for models is to predict the sentiment of movie reviews, negative, neutral, and positive, ranging from 0 to 1 without any prior examples (zero-shot learning), evaluating the capacity of models for natural language understanding and emotional tone assessment.
TagMyNews (TMN): Using the TagMyNews Dataset, we randomly picked 200 articles with seven categories, which are sourced from RSS feeds of popular newspapers (Vitale et al., 2012). The task aims to assess the efficacy of the model in accurately categorizing news articles based on their textual content utilizing a one-shot learning framework. The expected output schema is similar to the problem classification task.
LLaMA and CodeLLaMA. LLaMA (Touvron et al., 2023) is known for its auto-regressive architecture aimed at chat applications. The parameters of the series of LLaMA models range from 7 billion to 70 billion and are learned from publicly available online data amounting to 2 trillion tokens. Besides, LLaMA models have been fine-tuned with a combination of supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF), focusing on aligning the models to human preferences for helpfulness and safety. CodeLLaMA (Roziere et al., 2023) is one variant of LLaMA that specifically designed for programming contexts. This adaptation involves further fine-tuning LLaMA on datasets specific to coding, thereby boosting their code-related performance capabilities. Our experiments are carried out using the meta-llama/Llama-2-7b-chat-hf and codellama/CodeLlama-7b-Instruct-hf models, available on the public Hugging Face platform.
GPT3.5 and GPT4 (Achiam et al., 2023). GPT3.5 is a fined-tuned version of GPT3 developed in January 2022. It is built on Transformer architecture while it emphasizes decoder-only mechanisms. GPT4, which is known as the state-of-the-art LLM, has a substantial leap in capability and complexity compared to GPT-3.5. Both GPT3.5 and GPT4 are closed-source LLMs, we employ gpt-3.5-turbo-0125 and gpt-4-0125-preview models in our experiments through API access, specifically within text chat application contexts.

### v2

# Fault Detection for Large Language Models: Benchmark and Enhancement Evaluation and Improvement of Fault Detection for Large Language Models

@@ Abstract.
Large language models (LLMs) have recently achieved significant success across various application domains, garnering substantial attention from different communities. Unfortunately, even for the best LLM, many faults still exist that LLM cannot properly predict. Such faults will harm the usability of LLMs in general and could introduce safety issues in reliability-critical systems such as autonomous driving systems. How to quickly reveal these faults in real-world datasets that LLM could face is important, but challenging. The major reason is that the ground truth is necessary but the data labeling process is heavy considering the time and human effort. To handle this problem, in the conventional deep learning testing field, test selection methods have been proposed for efficiently evaluating deep learning models by prioritizing faults. However, despite their importance, the usefulness of these methods on LLMs is unclear, and lack of exploration. In this paper, we conduct the first empirical study to investigate the effectiveness of existing fault detection methods for LLMs. Experimental results on four different tasks (including both code tasks and natural language processing tasks) and four LLMs (e.g., LLaMA3 and GPT4) demonstrated that simple methods such as Margin perform well on LLMs but there is still a big room for improvement. Based on the study, we further propose MuCS, a prompt Mutation-based prediction Confidence Smoothing framework to boost the fault detection capability of existing methods. Concretely, multiple prompt mutation techniques have been proposed to help collect more diverse outputs for confidence smoothing. The results show that our proposed framework significantly enhances existing methods with the improvement of test relative coverage by up to 70.53%.
@@ 1. Introduction
In recent years, the potential of large language models (LLMs) has brought hope to general artificial intelligence (AGI). LLMs achieved promising and even the best results compared to other conventional methods in many areas, such as question answering (Zhou et al., 2022), sentiment analysis (Zhang et al., 2023a), and code understanding (Ma et al., 2023) that make LLMs become the first choice of deep learning models in such tasks. Besides the most famous application ChatGPT 111OpenAI, ChatGPT, https://chat.openai.com/, multiple LLMs have been proposed, e.g., LLaMA (Touvron et al., 2023), Alpaca (Taori et al., 2023), and Cerebras-GPT (Dey et al., 2023). More interestingly, focusing on software engineering (SE) tasks, a series of code-related LLMs with good programming ability have been released to help the software development, such as StarCoder (Li et al., 2023a) and CodeLLaMA (Roziere et al., 2023). Learning to use LLMs is already a basic skill for us in daily life.
@@
Despite the success, the same as other types of deep learning models, LLMs also suffer from different problems that limit their practical usage. For example, LLMs sometimes generate unrelated outputs to the inputs or incorrect results misaligned with established world knowledge, which is known as the hallucination problem (Zhang et al., 2023b). Besides, recent work (Dong et al., 2023a) showed that LLMs can be easily fooled by adversarial attacks and are not robust. Furthermore, considering SE tasks, existing code-related LLMs cannot capture equivalent semantics when the given natural language prompt expresses the same meaning in different languages (Peng et al., 2024). Those limitations remind us that LLMs are not always trustworthy, and it is necessary to carefully evaluate LLMs and quickly reveal the unreliable outputs of LLMs in practical usage.
@@
The basic way to evaluate LLMs and reveal their weakness is to prepare test data as diverse as possible to assess the performance of LLMs accordingly. However, preparing such test data is heavy work due to the complicated process of data collection, data cleaning, and data labeling. Especially, data labeling requires human effort with domain knowledge which is the most challenging part. Moreover, some LLMs are closed-source which requires funding expenses, e.g., it tasks 0.03$ per 1,000 tokens of inputs for GPT4 using OpenAI API. Therefore, it is almost impossible to thoroughly evaluate LLMs on the target downstream task for people who have limited budgets. As a result, how to efficiently test LLMs and filter the mispredicted data (unreliable outputs) of LLMs with less effort becomes an urgent problem.
@@
To tackle the data labeling issue mentioned above, in the field of deep learning testing, multiple fault detection methods that quickly identify fault data (data the model has incorrect prediction) in the test set without using the labeling information (Hu et al., 2024) have been proposed. Based on the information required, existing methods can be divided into learning-based methods and output-based methods. Those methods have been widely studied in the classical deep learning models and have proven to be effective in saving the labeling budget during testing. However, such studied classical deep learning models are constructively different from LLMs, e.g., LLMs are pre-trained using a large amount of pre-training data with diverse tasks which makes their prediction confidence (most fault detection methods rely on) unclear given a specific downstream task. Therefore, as no studies have explored the effectiveness of fault detection methods for LLMs, it is unknown whether we can directly employ existing methods to evaluate LLMs and save our effort or not.
@@
To bridge this gap, in this paper, we conduct the first study to investigate the effectiveness of existing fault detection methods for LLMs. The major challenge in the study is that most fault detection methods are proposed for classification tasks and use the output probabilities for data prioritization, but LLMs normally output a sequence of tokens without such a clear output probability for the given task. To solve this, we design the prompt to guide LLMs to produce the output with their confidence. For datasets that have more than two classes, we add examples (few-shot) in each class in the prompt as guidance to ask LLMs to predict new data. In total, our study covers nine fault detection methods including both learning-based methods, e.g., TestRank (LI et al., 2021), and output-based methods, e.g., ATS (Gao et al., 2022). Based on the experiment results collected from four datasets (e.g., sentiment analysis and code clone detection) and four LLMs (e.g., LLaMA3 and GPT4). We found that 1) LLMs are not well-calibrated and overconfident in clone detection, problem classification, and news classification tasks, 2) simple methods such as Margin perform better compared to others, and 3) however, existing fault detection methods perform relatively pool on LLMs compared their performance on classical deep learning models. There is a need to propose methods to further enhance existing fault detection methods.
@@
To do so, inspired by (Li et al., 2023b; Wang et al., 2021) which utilize mutation testing techniques to help test deep learning models, we introduce MuCS, a prompt Mutation-based prediction Confidence Smoothing framework to enhance existing fault detection methods. Specifically, we mutate the prompt by transformation methods to generate a set of prompt mutants first. Here, both text augmentation methods and code refactoring methods have been considered in MuCS. Then, we collect the output probabilities of all mutants and compute the average prediction confidence. Finally, we use the averaged confidence to perform fault detection using current fault detection methods. We compare the effectiveness of fault detection methods with and without our confidence smoothing method and found that using MuCS significantly enhances the performance of existing methods by up to 70.53%.
@@
This is the first study that explores the effectiveness of fault detection methods on LLMs. We found that 1) LLMs are overconfident in clone detection, problem classification, and news classification tasks; 2) the prediction confidence of LLMs is concentrated in a few intervals; and 3) simple methods perform relatively better than others.
@@ 2.1. Fault Detection for Deep Neural Networks
The early method DeepGini (Feng et al., 2020) has been proposed by Feng et al. to reveal inputs that are more likely been mispredicted by the model. DeepGini (Feng et al., 2020) defined a new Gini score to measure the uncertainty of deep learning models on the inputs. Later on, inspired by the mutation testing in the conventional SE field, Wang et al. (Wang et al., 2021) proposed to mutate both input samples and deep learning models and identify faults by computing the killing score. Their evaluation on more than 20 tasks demonstrated that mutation testing is useful for helping find bugs in deep learning models. Furthermore, Gao et al. (Gao et al., 2022) proposed an adaptive test selection (ATS) method. ATS not only considers detecting faults but also tries to find diverse faults (diverse here means faults are from different categories) in deep learning models. More recently, Bao et al. (Bao et al., 2023) first empirically proved that simple methods (i.e., output probability-based methods) perform well on fault detection. Then, they proposed a new method to compute the uncertainty scores by averaging the basic score of the input and scores from the neighbors of this input. Besides, Ma et al. (Ma et al., 2021) conducted a comprehensive empirical study to explore the potential of test selection methods and active learning methods on fault detection. They found that MaxP is the best in terms of the correlation between the MaxP score and the misclassification. Besides the SE community, researchers from the ML community also contributed to this field. Dan et al. (Hendrycks and Gimpel, 2016) built the first benchmark for detecting misclassified data and found that just using the confidence score is already a strong baseline for this task. Li et al. (LI et al., 2021) proposed TestRank which uses graph neural networks to learn the difference between benign inputs and faults for fault detection.
Different from these works, our work is the first to explore the effectiveness of fault detection methods for large language models. We found that existing methods cannot detect faults in LLMs well and proposed the use of prompt mutation to further enhance these methods.
@@ 2.2. Large Language Model Testing
Large language models have been the most used deep learning models for most of the tasks. Comprehensively testing or evaluating the performance of LLMs has become one of the most important directions in almost all communities.
@@
More recently, Yuan et al. (Yuan et al., 2023) compared LLMs with conventional code models that are fine-tuned for specific code tasks. They found that in the zero-shot setting, instruction-tuned LLMs have competitive performance to fine-tuned code models. In the one-shot setting, the guidance by one-shot example sometimes harms the performance of LLMs. Du et al. (Du et al., 2024) argued that existing works that evaluate the effectiveness of LLMs for code mainly focus on simple code tasks (function-level code synthesis) and built the first class-level code generation benchmark for the LLM evaluation. Based on their benchmark, they found that existing LLMs have significantly lower performance on class-level code generation tasks compared to method-level code generation tasks. There is still a big room to improve the ability of LLMs for code generation.
Existing LLM testing works mainly focus on constructing new and challenging datasets (by manual data collection and adversarial data generation, etc.) to validate the capability of LLMs, which can be seen as test augmentation techniques. In contrast, our work targets test selection from the perspective of software testing for LLM testing.
@@ 3. Empirical Study
Our study focuses on the problem of fault detection of LLMs. We consider both learning-based fault detection methods and output-probability-based methods in the literature. Based on the previous definition (Hu et al., 2024), given a LLM , an unlabeled test set , and a labeling budget , fault detection is to select a subset of such that where are the labels corresponding to , and is the performance measurement function. We follow the guidance of previous works (Hu et al., 2024; Ma et al., 2021; Hu et al., 2023) to select nine fault detection methods in our study. We exclude neuron coverage-based methods and surprise adequacy methods in our study since the parameter size of LLMs is too big (>7B) and the coverage and adequacy score are difficult to compute.
@@ 3.1. Fault Detection Methods
In total, we collect nine fault detection methods. Let be a set of test inputs and be the corresponding label set, and denote a given test sample and its true label, respectively. Let be the likelihood of belonging to class produced by the model .
@@
Max Probability (MaxP) prioritizes data based on the maximum output probability, expressed as . Data samples with lower MaxP scores are more likely to be faults.
@@
3.2. Datasets and Models
| Task | Class Number | Model | Test Size | Accuracy |
| Clone Detection (CD) | 2 | DeepSeekCoder | 1000 | 69.4% |
| GPT3.5 | 200 | 75.5% | ||
| GPT4 | 200 | 77.0% | ||
| Problem Classification (PC) | 5 | DeepSeekCoder | 1000 | 42.6% |
| GPT3.5 | 200 | 36.0% | ||
| GPT4 | 200 | 100.0% | ||
| Sentiment Analysis (Sentiment) | 3 | LLaMA3 | 1000 | 68.8% |
| GPT3.5 | 150 | 82.0% | ||
| GPT4 | 150 | 90.0% | ||
| TagMyNews (TMN) | 7 | LLaMA3 | 1400 | 46.7% |
| GPT3.5 | 200 | 60.0% | ||
| GPT4 | 200 | 82.0% |
We consider four datasets including both natural language classification tasks and programming language classification tasks, and four types of LLMs covering both open-sourced models and closed-source models. Due to the high cost of assessing close-source LLMs, we prepare relatively more test data for open-source LLMs than close-source LLMs.
Clone Detection (CD): Utilizing the BigCloneBench (Svajlenko et al., 2014) dataset, we randomly select 1000 and 200 items to investigate the code clone detection ability of open-source LLMs and closed-source LLMs, respectively. Models are evaluated on their ability to predict the presence of code clones (a probability score from 0 to 1, where 0 indicates no clone and 1 indicates a clone) between code snippet pairs, a critical task for enhancing software maintenance and development practices. Note that, some fault detection methods that require more than two classes such as ATS cannot apply to clone detection tasks.
Problem Classification (PC): Based on the Java250 (Puri et al., 2021) dataset from the CodeNet Project, problem classification focuses on the classification of programming challenges. 1000 samples and 200 samples across 5 types of problems are randomly chosen for evaluating open-source and closed-source LLMs. This task measures the model’s proficiency in categorizing coding problems using a singular example for each category, requiring output as a probability distribution across classes that sum to 1. We employ the one-shot setting (one-shot examples are randomly selected from the original dataset) for this task due to the poor performance of zero-shot evaluation (nearly random guess).
Sentiment Analysis (Sentiment): For this task, 1000 samples and 150 samples are extracted from the Sentiment Analysis of IMDB Movie Reviews dataset 222Autolabel, https://github.com/refuel-ai/autolabel to evaluate open-source and closed-source LLMs, respectively. The challenge for models is to predict the sentiment of movie reviews, negative, neutral, and positive, ranging from 0 to 1, evaluating the capacity of models for natural language understanding and emotional tone assessment.
TagMyNews (TMN): Using the TagMyNews Dataset, we randomly picked 1400 and 200 articles with seven categories, which are sourced from RSS feeds of popular newspapers (Vitale et al., 2012) for open-source and closed-source LLMs, respectively. This task aims to assess the efficacy of the model in accurately categorizing news articles based on their textual content. Due to the same reason as the problem classification task, we employ the one-shot setting for this task.
LLaMA3 and DeepSeekCoder. LLaMA3 (Touvron et al., 2023) is known for its auto-regressive architecture aimed at chat applications. The parameters of the series of LLaMA3 models range from 8 billion to 70 billion and are learned from publicly available online data amounting to 2 trillion tokens. Besides, LLaMA3 models have been fine-tuned with a combination of supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF), focusing on aligning the models to human preferences for helpfulness and safety. DeepSeeker-Coder consists of a series of code language models, which achieve competitive performance on different code-related tasks (Guo et al., 2024). DeepSeek-Coder not only beats most open-sourced LLMs but also surpasses existing advancing close-source LLMs like Codex and GPT-3.5. We employ the instruct version (meta-llama/Meta-Llama-3-8B-Instruct and deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct) available on the public Hugging Face platform for our experiments.
GPT3.5 and GPT4 (Achiam et al., 2023). GPT3.5 is a fined-tuned version of GPT3 developed in January 2022. It is built on Transformer architecture while emphasizing decoder-only mechanisms. GPT4, which is known as the state-of-the-art LLM, has a substantial leap in capability and complexity compared to GPT-3.5. Both GPT3.5 and GPT4 are closed-source LLMs, we employ gpt-3.5-turbo-0125 and gpt-4-0125-preview models in our experiments through API access, specifically within text chat application contexts.
3.3. Prompt Design
The quality of input prompts highly affects the output of LLMs. To obtain the output probabilities of LLMs given the input, we add guidance in the prompt to ask LLMs to produce their prediction confidence of each category. Figure 1 presents the template example of the code clone detection task. For multiple class classification tasks such as problem classification, we use one-shot examples to lead LLMs to produce more accurate prediction confidence.
@@ 3.4. Research Question
Our study aims to explore the effectiveness of existing fault detection methods on LLMs. Concretely, we answer the following two research questions:
RQ1: How confident are LLMs in their prediction? Most (seven out of nine) of the fault detection methods are purely output uncertainty-based. Therefore, checking the prediction confidence (an important indicator to measure the output uncertainty) produced by LLMs before running fault detection methods is necessary.